. Find the general solution of the partial differential equation

6uxx −5uxy −6uyy =52e3x+2y
1
Expert's answer
2020-04-15T09:43:24-0400

The given equation can be written as

"(6D^{2}-5DD'-6D'^{2})u=52e^{3x+2y}", where "D \equiv \dfrac{\partial}{\partial x}, D' \equiv \dfrac{\partial}{\partial y} ."


The auxiliary equation is

"6m^{2}-5m-6=0" , solving we get

"m = -\dfrac{2}{3}, \dfrac{3}{2}"

"C.F=f_{1}(y-\frac{2x}{3})+f_{2}(y+\frac{3x}{2})"

"P.I = \dfrac{1}{6D^2-5DD'-6D'^{2}} 52e^{3x+2y}\\
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\text{(Replace $D$ by 3 and $D'$ by 2)}\\
= \dfrac{1}{6 \cdot 3^2-5\cdot 3\cdot 2-6\cdot2^{2}} 52e^{3x+2y}\\
= \dfrac{1}{0}52e^{3x+2y}~~\text{(Rule fails)}\\
~~~~~~~~~~~~~~~~~~~~~~~~~~\text{(Differentiating denominator with respect to $D$} \\
~~~~~~~~~~~~~~~~~~~~~~~~~~~~\text{and multiplying the numerator by $x$)}\\
=x \dfrac{1}{12 D-5D'} 52e^{3x+2y}\\
~~~~~~~~~~~~~~~~~~~~~~~~~~\text{(Replace $D$ by 3 and $D'$ by 2)}\\
=x\dfrac{1}{36-10}52e^{3x+2y}\\
P.I=2xe^{3x+2y}"


The general solution is

"u(x,y) = C.F+P.I = f_{1}(y-\frac{2x}{3})+f_{2}(y+\frac{3x}{2})+2xe^{3x+2y}"